OpenStudy (anonymous):

When charcoal reacts in the presence of oxygen, CO and CO_2 are produced according to the following chemical reactions. \[C(s) + \frac{1}{2} O_{2}(g) \rightarrow CO(g)\] \[C(s) + O_{2} (g) \rightarrow CO_{2}(g)\] What would be the total mass of gas produced when \(400\) g of charcoal is reacted, assuming equal amounts are consumed in each reaction.

OpenStudy (anonymous):

How am i able to solve this? Anyone able to give me a hint? What formula do i use?

OpenStudy (australopithecus):

well find the moles of carbon first

OpenStudy (anonymous):

mole of carbon is 12.01

OpenStudy (australopithecus):

now what is the limiting reagent in the initial reaction?

OpenStudy (anonymous):

what is the limiting reagent? Sorry my chemistry is really rusty..

OpenStudy (australopithecus):

In a chemical reaction, the limiting reagent, also known as the "limiting reactant", is the substance which is totally consumed when the chemical reaction is complete. The amount of product formed is limited by this reagent since the reaction cannot proceed further without it. The other reagents may be present in excess of the quantities required to react with the limiting reagent.

OpenStudy (anonymous):

how am i able to calculate the limiting reagent then?

OpenStudy (australopithecus):

https://en.wikipedia.org/wiki/Limiting_reagent

OpenStudy (australopithecus):

it has been awhile since I have had to determine the limiting reagent let me refresh my memory but otherwise this problem is easy for me to solve :)

OpenStudy (australopithecus):

nevermind the limiting reagent is the charcoal

OpenStudy (australopithecus):

as it is what is going to be used up first

OpenStudy (australopithecus):

so once you have the moles you need to determine the amount of product since it is 1 to 1 you will have the same amount of moles of charcoal as you will for CO

OpenStudy (australopithecus):

the amount of moles of CO will equal the amount of moles of CO2 as it is still a 1 to 1 reaction thus all you have to do is take the moles and use the formula moles = grams/molecular mass find the molecular mass of CO2 using the periodic table and solve for grams

OpenStudy (anonymous):

\[n=\frac{400}{44.01}\]

OpenStudy (australopithecus):

wait you are incorrect

OpenStudy (australopithecus):

let me work through this you gave me the molecular weight not the moles

OpenStudy (anonymous):

i know..since i got it wrong :/

OpenStudy (anonymous):

okay thanks

OpenStudy (australopithecus):

Write down the molecular mass molecular mass of carbon is 12.01g/mol molecular mass of oxygen is 15.9994g/mol Thus the molcular mass of O2 is 15.9994*2 = 31.9988g/mol Molecular mass of CO = 28.094 Molecular mass of CO2 = 12.01 + 15.9994*2 = 44.0088g/mol Now we use the formula MEMORIZE THIS! moles = grams/molecular mass so we need to find the moles of carbon we know that the charcoal weighs 400g thus we have 400g/12.01g/mol = 33.3056mol of carbon all the reactions are 1 to 1 thus we end up with (for every 1 mole of carbon you have 1 mol of product) so 33.3056mol = x/44.0088g/mol = 1,465.7g This is assuming that the amount of oxygen present is unlimited. If the amount of oxygen was limited to 400g than you would need to find the limiting reagent

OpenStudy (anonymous):

the answer is \(1.2\) kg..

OpenStudy (australopithecus):

Ok then I'm going to assume there are 400 grams of oxygen sorry the question isnt very clear

OpenStudy (australopithecus):

No I'm sorry but my answer has to be correct 400g/12.01g/mol = 33.30mol 33.30mol* 44.0088g/mol = 1466g of CO2 I will get someone to check it for me give me a second

OpenStudy (anonymous):

well the answer i have is \(1.2\)kg and its a multiple choice and the choices are 0.93, 1.5, 2.5 which does not much what you have..

OpenStudy (australopithecus):

wel 1.5 matches

OpenStudy (australopithecus):

if you round up

OpenStudy (anonymous):

but that is apparently not the right answer..

OpenStudy (australopithecus):

oh crud

OpenStudy (australopithecus):

oh I miss read the question equal amounts used in both reactions :L

OpenStudy (australopithecus):

200g in one reaction and 200g in the other

OpenStudy (australopithecus):

so 200g/12.01g/mol = 16.65mol 16.65mol*28.094 = 467.8g then we have 200g/12.01g/mol = 16.65mol 16.65mol*44g/mol =732.87g thus 467.8g + 732.87g = 1.2kg

OpenStudy (australopithecus):

there is your answer sorry for not understanding the question I have really bad insomnia lol

OpenStudy (anonymous):

yes im just looking over what you all solved, in essence when you calculate the moles of charcoal you halve that value and then recalculate to products because from stechiometry equal amount of charcoal burns to form CO and CO2...

OpenStudy (anonymous):

thank you!

OpenStudy (australopithecus):

man I feel like an idiot not comprehending the question and wasting my time ugh lol